# The Medium Is Not the Message: Deconfounding Text Embeddings via Linear Concept Erasure

### v2

@@ 3.1 Category-level Data
The erased concept is the source for each of the four settings (Table 1). When generating clusters, we follow a standard practice and apply $k$-means to the text embeddings for each document Zhang et al. (2022). [^5] We set $k$ = 21, the total number of categories in the data. Improvements are robust to different $k$, see Fig. 9 in appendix.
@@
Table 2: Cluster alignment metrics on the “category-level” Comparative Agendas Project datasets (§3.1), before and after linear concept erasure. Here, the erased concept is the source (top row). We set $k$ = 21, the total number of categories in the CAP datasets. Erasure substantially improves cluster alignment for every combination of sources across all embedding models. Underlined scores indicate the highest value in each column.
@@ 3.3 Embedding Models
Our experiments use ten embedding models of varying sizes and dimensionality (appendix Table 12). This set includes multilingual and monolingual variants, as well as models with instruction fine-tuning: MiniLM [^6] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2, GIST-small, GIST-base, GIST-large Solatorio (2024), multilingual E5-small, E5-base, E5-large Wang et al. (2024), all-mpnet-base-v2 Song et al. (2020), Nomic-v2 Nussbaum and Duderstadt (2025), and MXB-large Li and Li (2023); Lee et al. (2024).
@@ 4 Primary Results
We first discuss the results on the category-level datasets, then turn to the event-level. In brief, erasure improves embeddings across the board—over all models, metrics, and datasets we study.
@@
Table 3: Document similarity search results on paired “event-level” U.S. Supreme Court Summaries (3.2), before and after linear concept erasure. Here, the erased concept is the document’s source. Erasure improves recall for every setting and model. Underlined scores indicate the highest value in each column.
4.1 Category-level
In all four source pairings from the CAP dataset, erasing source-specific information with LEACE consistently improves clustering quality (Table 2). In the Bills–Newspapers comparison, all ten models show marked improvements, with gains in ARI ranging from +0.104 (E5-large) to +0.157 (GIST-small), and purity increases as high as +0.169 (GIST-small). Although the magnitude of improvement varies, this pattern persists in the Orders–Newspapers comparison. While most models benefit substantially, multilingual models such as E5-small and E5-large show only marginal gains, suggesting that source signal may be less distinct in this pairing.
@@ 4.2 Event-Level
At the event level, we present the results with Recall@10 and Recall@1, because only one document is deemed relevant for each query [^7] In Appendix F, we summarized the results of a qualitative analysis using pairs of legal case summaries sourced from Wikipedia and LexisNexis, using multilingual E5 as the embedding model..
@@
Applying LEACE consistently improves retrieval performance on the SCOTUS summary data (Table 10). In both Wikipedia pairings, improvements are large and especially pronounced for the E5 family. For instance, on LexisNexis-Wikipedia, E5-small gains +0.177 in Recall@1 and E5-base +0.153.
@@
Turning now to multilingual data, we observe that LEACE can be extremely effective, even with already-multilingual embeddings (Table 4).
@@
Table 4: Document similarity search results on paired “event-level” multilingual Swiss Court Case Summaries (3.2), before and after linear concept erasure. Here, the concept is the document’s language. Once again, erasure improves recall of the paired item in all cases, in some instances improving smaller models over their larger counterparts. Underlined scores indicate the highest value in each column.
SemEval News Articles
To avoid bludgeoning the reader with positive results, we briefly outline the results on our other multilingual dataset: all ten models again benefit from erasure (Table 6 in the appendix).
@@ 5 Erasure helps, but can it hurt?
The results from the previous section appear conclusive: linear concept erasure removes spurious information from embeddings that distort similarities. At the same time, we might wonder: is it possible that erasure degrades the embeddings in subtle ways that our evaluations fail to detect? Although LEACE is designed to avoid unwanted distortions, it is possible the trained eraser removes “desirable” information that may support other tasks.
In this section, we explore this question with additional evaluations on out-of-distribution (OOD) benchmarks. The experiments are designed to answer whether applying an eraser trained for a specific domain unintentionally harms general-purpose semantic representations. While our main experiments in the previous section target domain-specific differences, real-world deployment of embedding models often involves cross-domain tasks. We thus benchmark our models against diverse evaluation datasets from MTEB, Muennighoff et al. (2023), determining whether erasers trained to isolate certain information also degrade performance in unrelated tasks.
@@ 5.2 MTEB Results
On both the legal and news retrieval tasks, the trained erasers do not harm performance (as measured by the average NDCG@10). See Fig. 3 for legal retrieval; per-task performance (Fig. 5) and news retrieval (Fig. 6) are in the appendix. Given the domain overlap, we had hypothesized that the Legal eraser might improve legal retrieval somewhat, but only one task sees a marginal improvement (AILACasedocs), from 0.197 to 0.218 (Table 7 in appendix). That said, the results are still positive overall, indicating that both the CAP and Legal erasers operate robustly in OOD retrieval tasks with both small and large models.
@@
We evaluate the four model–eraser combinations on eight well-established STS benchmarks, covering both monolingual and crosslingual settings in the news domain (Fig. 7 and Table 9 in the appendix). The evaluation metric is the Spearman correlation between embedding cosine similarities and ground-truth semantic similarity. LEACE does not degrade performance over tasks, with most scores either unchanged or showing negligible increases.
@@ 6 Additional Findings
Taking the embeddings of the German-French Swiss court summaries, the first principal component (PC1) forms two clearly separable clusters corresponding directly to the text’s language (Fig. 2). After applying LEACE, the clusters collapse into a single, overlapping distribution, an indication that language identity is no longer linearly separable in the embedding space.
@@
We apply PCA to the original embeddings from each event-level dataset (SCOTUS, SemEval, Swiss Court Cases) and record the proportion of total variance explained by PC1. A high proportion of explained variance suggests that PC1 encodes a dominant direction in the embedding space, which will tend to correspond to the concept targeted by LEACE (i.e., the source or language, per Fig. 2).
Fig. 4 shows a strong positive correlation ($r=0.79$, $p<0.001$) between the proportion of variance explained by PC1 and the percentage improvement in Recall@1. This result indicates that LEACE is more effective when the removable concept aligns with dominant directions in the embedding space.
Given the above findings, why not use PCA to perform erasure instead, along the lines of Bolukbasi et al. (2016)? We find positive but less consistent results than LEACE on our tasks, and strongly degraded MTEB performance (Appendix E).
@@
Erasure improves already multilingual models on with multilingual tasks, so can it help with bitext mining—retrieving translation pairs via similarity search? Improvements are not as universally strong, but we do observe state-of-the-art results on a few leaderboard tasks from Enevoldsen et al. (2025), and erasure never harms performance (details in Appendix A).
@@ 7 Conclusion
For applied practitioners working with large text collections from multiple sources or languages—a regular occurrence—our results tell a fairly unambiguous story: apply linear erasure to any document embeddings before working with them to remove confounding information. While there are cases where it may not work, it does not seem to damage representations (see below), and comes at a minimal computational cost.
8 Limitations
The primary limitation of our method is its dependence on per-document metadata or labels. If some undesirable low-level pattern in the data distribution is suspected but not known—say, an unreported change in how a corpus was collected over a long time period—then a user will have to first apply some possibly-unsupervised labeling method.
Another shortcoming arises when metadata is available but the categories are too numerous relative to the total number of items. For instance, the paired within-language (en–en) SemEval Task 8 news articles come from dozens of sources, with many sources only being represented by a handful of articles. In contrast to removing the language in the multilingual data (Table 6), removing the source label does not improve retrieval results over the baseline. A possible direction for future work is to first combine similar sources into larger categories (e.g., local vs. national newspapers), then erase the category label.
A final limitation was first noted by Huang et al. (2024). They use LEACE as a baseline in multilingual retrieval contexts, removing language information as we do, but find mixed results. Hence, LEACE may not always help in all contexts. One initial hypothesis is that our tasks, while realistic, differ from the standard benchmark data that models are trained on, which may lead to saturated in-domain performance that does not transfer out-of-domain. It may also be that retrieval setups, with the distinct (short query, document) rather than our (document, document) structure, have characteristics that make them less amenable to erasure. In future work, we plan to explore these hypotheses and find an explanation for such discrepancies.
@@ Appendix A Bitext Mining Results
In several cases, there is a marked increase, yielding state-of-the-art scores on three tasks that appear on the public leaderboard (even with a different base model class, Table 5 in appendix). Generally, though, the improvements are much smaller than those in our main experiments, with over half of the 28 tasks showing less than a 0.01 change (although no tasks decrease more than $-0.01$). First applying LEACE is therefore a simple step when bitext mining; even if it may not always help, it is unlikely to hurt.
@@ Appendix B SemEval English & Non-English News Results
The results on testing LEACE on the SemEval 2022 Task 8 dataset are presented in Table 6. All models benefit from LEACE, with consistent improvements in both Recall@10 and Recall@1. The E5-small model shows the strongest gains overall: +0.202 (Recall@10) and +0.236 (Recall@1). High-performing large models like E5-large and MXB-large achieve further enhancements of up to +0.156 in Recall@1. Smaller models also gain notable increases. For instance, MiniLM gains +0.183 (Recall@10) and +0.127 (Recall@1), respectively. These improvements highlight LEACE’s utility in reducing source bias and improving semantic alignment in multilingual event representations. Nomic-v2, which already has high scores before LEACE, showed modest increases, likely due to saturation. In general, LEACE proves effective even under high-resource, multilingual scenarios.
@@ Appendix C MTEB Evaluation Results
We report the full evaluation results of the CAP and Legal erasers on three MTEB benchmark groups: Legal Retrieval, News Retrieval, and STS News Tasks. Each setting involves comparing model performance before and after LEACE-based erasure, across two embedding models (MiniLM and E5-base), as shown in Table 7, Table 8, and Table 9 and Fig. 5, Fig. 6 and Fig. 7.
@@ Appendix E Additional PCA Analysis
We create a baseline by removing PC1 from the embedding space, and evaluate it in the event-level setting using the SCOTUS dataset (Table 10). Overall, the baseline occasionally helps and can even marginally outperform LEACE in a few cases, but its effectiveness appears unstable, heavily dependent on the particular setting and model used (although it is effective for the E5 family for most configurations). Furthermore, in some cases, it performs worse than applying no erasure at all. There is also a final catch: removing the learned PC1 from OOD embeddings does dramatically degrade performance on MTEB tasks (Table 11), unlike LEACE (Fig. 8).
@@ E.1 Event-Level Results on SCOTUS Case Summaries
Table 10 reveals the results of applying the baseline, which removes the first principal component (PC1) from the embedding space, in the event-level setting on the SCOTUS dataset. While it sometimes improves over the original embeddings and occasionally outperforms LEACE (especially for the E5 family), its performance is inconsistent across models and configurations, and it can underperform even relative to no erasure.
@@ E.2 MTEB Evaluation Results
Table 11 shows the results of applying the baselines, derived from both CAP and SCOTUS datasets, on the MTEB legal retrieval tasks. In all cases, this PC1 removal leads to a drastic performance drop for both MiniLM and E5-base models. As observed in the comparison between the two approaches in Fig. 8, in contrast, LEACE erasures maintain retrieval quality, highlighting its robustness.
@@
Appendix F Qualitative Analysis
Our examination of legal case summaries provides insights into these dynamics. Summaries that experience the largest changes after applying LEACE are generally shorter, focus primarily on legal rulings, and lack factual idiosyncrasies. Without LEACE-based deconfounding, these summaries present substantial challenges for similarity-based text retrieval because of their limited semantic overlap regarding specific facts and rules (see also, e.g. Fan et al. 2025). For example:
@@ Appendix F Qualitative Analysis
On the other hand, summaries that display minimal change tend to incorporate both factual details and judgments, offer broader contextual framing, discuss subsequent impacts, and frequently integrate direct quotations from court decisions. For instance:
@@
These findings underscore LEACE’s distinctive advantage in scenarios where domain experts can clearly identify and specify features irrelevant to downstream similarity tasks, highlighting its potential value within human-in-the-loop frameworks that leverage expert knowledge to detect and eliminate confounding factors.

### v3

@@ 3.1 Category-level Data
The erased concept is the source for each of the four settings (Table˜1). When generating clusters, we follow a standard practice and apply $k$-means to the text embeddings for each document Zhang et al. (2022). [^5] We set $k$ = 21, the total number of categories in the data. Improvements are robust to different $k$, see Fig. 10 in appendix.
@@
Table 2: Cluster alignment metrics on the “category-level” Comparative Agendas Project datasets (§3.1), before and after linear concept erasure. Here, the erased concept is the source (top row). We set $k$ = 21, the total number of categories in the CAP datasets. Erasure substantially improves cluster alignment for every combination of sources across all embedding models. Bolded scores indicate performance improvements after erasure; underlined scores mark the highest value in each column.
@@ 3.3 Embedding Models
Our experiments use ten embedding models of varying sizes and dimensionality (appendix Table˜12). This set includes multilingual and monolingual variants, as well as models with instruction fine-tuning: MiniLM [^6] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2, GIST-small, GIST-base, GIST-large Solatorio (2024), multilingual E5-small, E5-base, E5-large Wang et al. (2024), all-mpnet-base-v2 Song et al. (2020), Nomic-v2 Nussbaum and Duderstadt (2025), and MXB-large Li and Li (2023); Lee et al. (2024).
@@ 4 Primary Results
We first discuss the results on the category-level datasets, then turn to the event-level. In brief, erasure improves embeddings across the board—over all models, metrics, and datasets (when performance is not already saturated). A summary of results for a single model is in Fig. 2.
4.1 Category-level
In all four source pairings from the CAP dataset, erasing source-specific information with LEACE consistently improves clustering quality (Table˜2). In the Bills–Newspapers comparison, all ten models show marked improvements, with gains in ARI ranging from +0.104 (E5-large) to +0.157 (GIST-small), and purity increases as high as +0.169 (GIST-small). Although the magnitude of improvement varies, this pattern persists in the Orders–Newspapers comparison. While most models benefit substantially, multilingual models such as E5-small and E5-large show only marginal gains, suggesting that source signal may be less distinct in this pairing.
@@ 4.1 Category-level
Table 3: Document similarity search results on paired “event-level” U.S. Supreme Court Summaries (3.2), before and after linear concept erasure. Here, the erased concept is the document’s source. Erasure improves recall for every setting and model.
@@ 4.2 Event-Level
At the event level, we present the results with Recall@10 and Recall@1, because only one document is deemed relevant for each query.
@@
Applying LEACE consistently improves retrieval performance on the SCOTUS summary data (Table˜10). In both Wikipedia pairings, improvements are large and especially pronounced for the E5 family. For instance, on LexisNexis-Wikipedia, E5-small gains +0.177 in Recall@1 and E5-base +0.153.
@@
Turning now to multilingual data, we observe that LEACE can be extremely effective, even with already-multilingual embeddings (Table˜4).
SemEval News Articles
To avoid bludgeoning the reader with positive results, we briefly outline the results on our other multilingual dataset: all ten models again benefit from erasure (Table˜6 in the appendix).
@@
Table 4: Document similarity search results on paired “event-level” multilingual Swiss Court Case Summaries (3.2), before and after linear concept erasure. Here, the concept is the document’s language. Once again, erasure significantly improves recall of the paired item in all cases. The only exception is one instance where retrieval performance is already perfect before erasure. In some cases, erasure even allows smaller models to outperform their larger counterparts.
4.2.1 Qualitative Analysis
Our examination of legal case summaries provides insights into these dynamics. Summaries that experience the largest changes after applying LEACE are generally shorter, focus primarily on legal rulings, and lack factual idiosyncrasies. Without LEACE-based deconfounding, these summaries present substantial challenges for similarity-based text retrieval because of their limited semantic overlap regarding specific facts and rules (see also Fan et al. 2025). We provide examples in Section˜F.1.
On the other hand, summaries that display minimal change tend to incorporate both factual details and judgments, offer broader contextual framing, discuss subsequent impacts, and frequently integrate direct quotations from court decisions. An example can be found in Section˜F.2.
These findings underscore LEACE’s distinctive advantage in scenarios where domain experts can clearly identify and specify features irrelevant to downstream similarity tasks, highlighting its potential value within human-in-the-loop frameworks that leverage expert knowledge to detect and eliminate confounding factors. They also show LEACE is less useful when relevance relies on higher-order discourse features such as blending facts, judgments, and framing, which it cannot account for.
@@ 5 Erasure helps, but can it hurt?
The results from the previous section appear conclusive: linear concept erasure effectively removes spurious information from embeddings that distorts similarities. At the same time, we must ask whether erasure might also degrade embeddings in subtle ways that our evaluations fail to detect. Although LEACE is designed to minimize unwanted distortions, the trained eraser may inadvertently remove “desirable” information that may support other tasks.
In this section, we address this question through additional evaluations on out-of-distribution (OOD) benchmarks. These experiments test whether applying an eraser trained for a specific domain unintentionally harms general-purpose semantic representations. While our main experiments in the previous section focused on domain-specific differences, real-world deployment of embedding models often requires robust cross-domain performance. We thus benchmark our models against diverse evaluation datasets from MTEB Muennighoff et al. (2023) to assess whether erasers trained to isolate certain information also degrade performance in unrelated tasks.
@@ 5.2 MTEB Results
On both the legal and news retrieval tasks, the trained erasers do not harm performance (as measured by the average NDCG@10). See Fig.˜4 for legal retrieval; per-task performance (Fig.˜6) and news retrieval (Fig.˜7) are in the appendix. Given the domain overlap, we had hypothesized that the Legal eraser might improve legal retrieval somewhat, but only one task sees a marginal improvement (AILACasedocs), from 0.197 to 0.218 (Table˜7 in appendix). That said, the results are still positive overall, indicating that both the CAP and Legal erasers operate robustly in OOD retrieval tasks with both small and large models.
@@
We evaluate the four model–eraser combinations on eight well-established STS benchmarks, covering both monolingual and crosslingual settings in the news domain (Fig.˜8 and Table˜9 in the appendix). The evaluation metric is the Spearman correlation between embedding cosine similarities and ground-truth semantic similarity. LEACE does not degrade performance over tasks, with most scores either unchanged or showing negligible increases.
@@ 6 Additional Findings
Taking the embeddings of the German-French Swiss court summaries, the first principal component (PC1) forms two clearly separable clusters corresponding directly to the text’s language (Fig.˜3). After applying LEACE, the clusters collapse into a single, overlapping distribution, an indication that language identity is no longer linearly separable in the embedding space.
@@
To test this, we apply PCA to the original embeddings from each event-level dataset (SCOTUS, SemEval, Swiss Court Cases) and record the proportion of total variance explained by PC1. A high proportion of explained variance suggests that PC1 encodes a dominant direction in the embedding space, which will tend to correspond to the concept targeted by LEACE (i.e., the source or language, per Fig.˜3). Indeed, Fig.˜5 shows a strong positive correlation ($r=0.79$, $p<0.001$) between the proportion of variance explained by PC1 and the percentage improvement in Recall@1. This result indicates that LEACE is more effective when the removable concept aligns with dominant directions in the embedding space.
Given these findings, one might ask why not use PCA for erasure instead, following Bolukbasi et al. (2016). We observe positive but less consistent results than LEACE on our tasks, along with a strong degradation in MTEB performance (Appendix˜E).
@@
Erasure improves already multilingual models on with multilingual tasks, so can it help with bitext mining—retrieving translation pairs via similarity search? Our further experiments show that improvements are not uniformly strong, but we do achieve state-of-the-art results on a few leaderboard tasks from Enevoldsen et al. (2025), and erasure never reduces performance (details in Appendix˜A).
@@ 7 Conclusion
For applied practitioners working with large text collections from multiple sources or languages—a common scenario—our results offer a clear recommendation: apply linear erasure to document embeddings before use to remove confounding information. While there may be cases where it is less effective, the method does not appear to harm representations (see below) and incurs only minimal computational cost.
Limitations
The primary limitation of our method is its dependence on per-document metadata or labels. If an undesirable low-level pattern in the data distribution is suspected but not known—say, an unreported change in how a corpus was collected over a long time period—then the user must first apply some possibly-unsupervised labeling method. Although confounder labels are available for many tasks, reliance on such labels constrains the broader applicability of our proposed methods. We explored approaches that automatically generate features; however, these did not lead to measurable improvements in downstream retrieval or similarity comparison tasks. Other works have developed unsupervised techniques to debias or erase neural representations Kim et al. (2019); Seo et al. (2022); Yang et al. (2025). We leave a deeper exploration of such methods in our context for future work. One direction is to integrate sparse autoencoders (e.g., Movva et al., 2025; Paulo et al., 2025) that extract features whose utility can be interpreted and validated by human experts.
Another shortcoming arises when metadata is available but the categories are too numerous relative to the total number of items. For example, in the paired within-language (en–en) SemEval Task 8 news articles, the data originate from dozens of sources, many of which are represented by only a handful of articles. In contrast to removing the language label in the multilingual data (Table˜6), removing the source label does not improve retrieval results over the baseline. A possible direction for future work is to first merge similar sources into broader categories (e.g., local vs. national newspapers) before applying label erasure.
A final limitation was first noted by Huang et al. (2024), who used LEACE as a baseline in multilingual retrieval contexts. While they similarly removed language information, their results were mixed, suggesting that LEACE may not be effective in all settings. One hypothesis is that our tasks, though realistic, differ from the standard benchmark data on which models are typically trained, leading to saturated in-domain performance that does not transfer well out-of-domain. Another possibility is that retrieval setups—with their distinct (short query, document) structure, as opposed to our (document, document) structure—may be less amenable to erasure. We plan to explore these hypotheses to help explain such discrepancies.
Acknowledgments
Funding provided in part by the ETH AI Center and Innosuisse (project number: 122.540 IP-ICT). Thanks to Maria Antoniak for pointing out the connection to Authorless Topic Models, and to Jacob Conway for his insightful suggestions on task transferability.
@@ Appendix A Bitext Mining Results
In several cases, there is a marked increase, yielding state-of-the-art scores on three tasks that appear on the public leaderboard (even with a different base model class, Table˜5 in appendix). Generally, though, the improvements are much smaller than those in our main experiments, with over half of the 28 tasks showing less than a 0.01 change (although no tasks decrease more than 0.01-0.01). First applying LEACE is therefore a simple step when bitext mining; even if it may not always help, it is unlikely to hurt.
@@ Appendix B SemEval English & Non-English News Results
The results on testing LEACE on the SemEval 2022 Task 8 dataset are presented in Table˜6. All models benefit from LEACE, with consistent improvements in both Recall@10 and Recall@1. The E5-small model shows the strongest gains overall: +0.202 (Recall@10) and +0.236 (Recall@1). High-performing large models like E5-large and MXB-large achieve further enhancements of up to +0.156 in Recall@1. Smaller models also gain notable increases. For instance, MiniLM gains +0.183 (Recall@10) and +0.127 (Recall@1), respectively. These improvements highlight LEACE’s utility in reducing source bias and improving semantic alignment in multilingual event representations. Nomic-v2, which already has high scores before LEACE, showed modest increases, likely due to saturation. In general, LEACE proves effective even under high-resource, multilingual scenarios.
@@ Appendix C MTEB Evaluation Results
We report the full evaluation results of the CAP and Legal erasers on three MTEB benchmark groups: Legal Retrieval, News Retrieval, and STS News Tasks. Each setting involves comparing model performance before and after LEACE-based erasure, across two embedding models (MiniLM and E5-base), as shown in Table˜7, Table˜8, and Table˜9 and Fig.˜6, Fig.˜7 and Fig.˜8.
@@ Appendix E Additional PCA Analysis
We create a baseline by removing PC1 from the embedding space, and evaluate it in the event-level setting using the SCOTUS dataset (Table˜10). Overall, the baseline occasionally helps and can even marginally outperform LEACE in a few cases, but its effectiveness appears unstable, heavily dependent on the particular setting and model used (although it is effective for the E5 family for most configurations). Furthermore, in some cases, it performs worse than applying no erasure at all. There is also a final catch: removing the learned PC1 from OOD embeddings does dramatically degrade performance on MTEB tasks (Table˜11), unlike LEACE (Fig.˜9).
@@ E.1 Event-Level Results on SCOTUS Case Summaries
Table˜10 reveals the results of applying the baseline, which removes the first principal component (PC1) from the embedding space, in the event-level setting on the SCOTUS dataset. While it sometimes improves over the original embeddings and occasionally outperforms LEACE (especially for the E5 family), its performance is inconsistent across models and configurations, and it can underperform even relative to no erasure.
@@ E.2 MTEB Evaluation Results
Table˜11 shows the results of applying the baselines, derived from both CAP and SCOTUS datasets, on the MTEB legal retrieval tasks. In all cases, this PC1 removal leads to a drastic performance drop for both MiniLM and E5-base models. As observed in the comparison between the two approaches in Fig.˜9, in contrast, LEACE erasures maintain retrieval quality, highlighting its robustness.
@@
Appendix F Examples for Qualitative Analysis
We provide here examples for legal case summaries that experience the largest and least changes after applying LEACE.
F.1 Examples with The Largest Changes
@@
F.2 An Example with The Least Changes